1 a.     Consider a baker selling cakes. There are 10 customers, 3 are willing to pay up to £5 for cakes, 4 are willing to pay up to £8, and 3 are willing to pay up to £10. 

i.If the baker can price discriminate, and costs are zero, what are the highest profits that he can earn. You must show your calculations clearly (note in this case, profits are equal to revenues because costs are zero). ii. If the baker cannot price discriminate, and costs are zero, what is the price she should sell the cakes at to maximise her profits? You must show your calculations clearly (note in this case, profits are equal to revenues because costs are zero ) iii.         Explain clearly whether, and if so how, taking into consideration the costs of baking each cake might affect your answers to “i" and “ii”. 



1
Expert's answer
2021-03-29T10:09:41-0400

i)

"5\times3+8\times4+10\times3=77"

ii)

"\frac{77}{10}=7.7"

iii) "\pi=TR-TC"